OpenStudy (anonymous):

56 Prime Composite:1,2,4,7,8,14,28,56 Composite:1,2,4,6,7,8,9,14,28,54 Composite:1.2.14.56 65 Prime Composite:1,65 Composite:1,3,5,13,21,65 Composite:1,5,13,65

OpenStudy (dayakar):

56 is composite number

OpenStudy (anonymous):

which compoiste

OpenStudy (anonymous):

and what about 65

OpenStudy (dayakar):

1,2,4,7,8,14,28,56

OpenStudy (anonymous):

whats 65

OpenStudy (dayakar):

do u know how it will com

OpenStudy (anonymous):

what the answer to 65

OpenStudy (dayakar):

65 = 1*65 = 5*13 composite 1,5,13,65

OpenStudy (dayakar):

r u clear

OpenStudy (anonymous):

okay but type in another question brb

OpenStudy (dayakar):

what

OpenStudy (anonymous):

go to my otheer question its at the top